1. Find the slope of the line that passes through the points (-3, -5) and (3, 2).

2. Find the equation, in slope-intercept form, of the line that passes through the points (-3, 1) and (-2, -5).

3. Find the equation, in standard form, with all integer coefficients, of the line perpendicular to x – 2y = 24 and passing through (1, 3).

4. Solve the system of equations using the substitution method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
7x + y = 1
3x + 2y = -9

Answer by tutorcecilia(2105)

1. Find the slope of the line that passes through the points (-3, -5) and (3, 2)
m=(y2-y1)/(x2-x1) [formula for the slope of a line]
m=(2-(-5))/(3-(-3)) [plug-in the values]
m=(2+5)/(3+3)=7/6